OpenStudy (anonymous):

The denominator of the fraction 4/15 is increased by an amount so that the value of the resulting fraction is 1/5. Determine the amount by which the denominator was increased.

OpenStudy (anonymous):

Do you have an idea of how to write the equation?

OpenStudy (anonymous):

A little number sense will find it too. 4/x = 1/5?

OpenStudy (anonymous):

then what do i do?

OpenStudy (anonymous):

Think about it.

OpenStudy (anonymous):

It's like reducing fractions to lowest terms, but in reverse because you are finding higher terms.

OpenStudy (anonymous):

The actual equation to find what the question is asking is\[\frac{4}{15+x}=\frac{1}{5}\] You can cross-multiply and all that to solve the proportion, but a little number sense will find the answer much faster.

OpenStudy (anonymous):

i got 5